### v2

# ChatModel: Automating Reference Model Design and Verification with LLMs

Abstract
As the complexity of integrated circuit designs continues to escalate, the functional verification becomes increasingly challenging. Reference models, critical for accelerating the verification process, are themselves becoming more intricate and time-consuming to develop. Despite the promise shown by large language models (LLMs) in code programming, effectively generating complex reference models remains a significant hurdle. To address these challenges, we introduce ChatModel, the first LLM-aided agile reference model generation and verification platform. ChatModel streamlines the transition from design specifications to fully functional reference models by integrating design standardization and hierarchical agile modeling. Employing a building-block generation strategy, it not only enhances the design capabilities of LLMs for reference models but also significantly boosts verification efficiency. We evaluated ChatModel on 300 designs of varying complexity, demonstrating substantial improvements in both efficiency and quality of reference model generation. ChatModel achieved a peak performance improvement of 55.02% compared to alternative methods, with notable enhancements in generation stability, and delivered a 9.18× increase in its capacity to produce reference model designs. Furthermore, it accelerated the iterative process of reference model design and validation by an average of 5.90× compared to traditional approaches. These results highlight the potential of ChatModel to significantly advance the automation of reference model generation and validation.
1 Introduction
The rapidly increasing complexity of integrated circuit (IC) designs has significantly prolonged functional verification cycles, posing a critical bottleneck in agile hardware development [37, 21, 32]. In this context, model-driven verification (MDV) is recognized as a highly effective methodology and has been widely applied in preliminary chip designs, such as neural processors and artificial intelligence accelerators [66, 14, 52, 43, 22, 33]. The models in MDV are typically written in high-level abstraction languages (i.e. Python [29, 3, 48], C++ [70, 67], MATLAB [19, 74], or SystemC [7, 1, 50, 2]) by engineers and serve as golden references for hardware designs. As shown in Figure 2, by systematically comparing the model outputs with those of the Device Under Test (DUT) under identical input stimuli, developers can effectively identify design flaws and accelerate the debugging and iteration process [53].
In the MDV framework, open-source simulators or loosely-timed modeling approaches are often employed to rapidly develop models [6, 23, 8, 5]. However, the execution processes within these models exhibit substantial discrepancies compared to actual hardware execution, thereby restricting their effectiveness in diagnosing errors within the DUT. For the development of specialized designs, developers must consider timing factors to construct more accurate models. However, building high-fidelity reference models requires a deep understanding of the behavior and verification requirements of the target hardware. As hardware designs grow increasingly intricate, the expertise and effort required expand exponentially, hindering agile model development.
With the advent of large language models (LLMs), a revolutionary paradigm shift has occurred in the field of natural language processing (NLP) [47]. Recent studies have showcased their impressive capabilities in code generation for programming languages [85, 56, 28, 44, 77, 90] such as C++ and Python, as well as hardware description languages [79, 16, 78, 75] such as Verilog. However, their application in constructing reference models remains underexplored. General-purpose LLMs encounter numerous challenges when applied to reference modeling (detailed further in Appendix A), including context and token limitations, hallucinations, and limited self-correction capabilities. These issues undermine the quality and stability of the outputs, restricting LLMs to small-scale designs. Moreover, the increasing diversity of design specifications and hardware architectures further exacerbates these issues, making it challenging for LLMs to meet the demands of constructing reference models.
To overcome these limitations, we introduce ChatModel, an end-to-end platform designed for agile generation and verification of reference models, as illustrated in Figure 2. The ChatModel platform leverages a multi-agent LLM system comprising two integrated groups: one standardizes design specifications, while the other automates the generation and validation of reference models. By adaptively decomposing complex design tasks into specialized subtasks, ChatModel facilitates the rapid construction of corresponding design Intermediate Representations (IR) and their associated reference models. Consequently, the ChatModel enables agile reference modeling, thereby substantially reducing both the time and effort required for the model design and validation. The contributions of our work are summarized as follows: In the first phase, specification standardization agents group automates the standardization process based on design specifications, transforming non-standard specifications into a standardized Intermediate Representation (IR). This IR functions as a novel domain-specific language for describing reference models. In the second phase, reference model generation agents group utilizes the generated Design IR to implement the Hierarchical Agile Modeling (HAM) flow for reference model generation. HAM employs an adaptive task planning algorithm to devise the construction strategy for the reference model, facilitating automated generation and verification from module to subsystem to system in a “building block” manner.
Figure 1: Model-Driven Verification Flow
Figure 2: The Workflow of ChatModel
  • We propose the ChatModel, the first multi-agent LLM-based platform for automated design and verification of reference models, enabling agile hardware verification.
  • We propose the Design IR, a novel domain-specific language specifically designed for LLM to accurately generate reference models, which effectively transforms complex specifications into a clear directed acyclic graph.
  • We introduce a Hierarchical Agile Modeling (HAM) flow that realizes agile generation and verification of models through a modular generation methodology.
  • We develop a benchmark, ModelEval, to comprehensively evaluate the capabilities of LLM in the generation of complex reference models.
  • Our comprehensive experiments demonstrate that ChatModel outperforms other benchmark methods, offering enhanced generative capabilities and more stable generation scales. Compared to general LLMs, it achieved a 55.02% improvement in functional pass rate and enhanced the LLM reference model generation ability by 9.18x. Furthermore, it delivers an average 5.90× speedup in model generation compared to experienced engineers.
The remainder of the paper is organized as follows: Section 2 discusses related work. Section 3 introduces the ChatModel platform. Sections 4 &  5 introduce the two phases of the ChatModel and some key components. Section 6 evaluates ChatModel and analyzes the results. Finally, Section 7 concludes this paper.
2 Related Works
Traditional Reference Model Design: As a vital component in the verification process, reference models are typically developed using programming languages at varying abstraction levels, tailored to the functionality of the DUT [65, 54]. For image or signal processing designs, Python [68, 62, 4] and MATLAB [38, 91] are commonly employed for modeling, while C++ [26, 51, 67] is predominantly used for constructing reference models for logic functions and CPU designs. Recently, SystemC with Transaction-Level Modeling 2.0 (TLM 2.0) has achieved widespread adoption for reference model design, attributable to its capacity to facilitate modeling across multiple levels of abstraction [25, 31, 59, 39, 17]. For larger System-on-Chip (SoC) designs, simulators like gem5 [9, 63, 72] or SST [69] are employed for model development to verify hardware system correctness, but their limited configurability results in coarse simulation granularity. Although these approaches have contributed to the advancement of reference model development, they require developers to possess substantial expertise in hardware and software programming, often necessitating a considerable and sustained investment of resources and time.
LLM in Code Generation: With the emergence of models such as ChatGPT [35] and Claude [10], LLMs have demonstrated their powerful capabilities in automating code generation from natural language descriptions [15, 88, 92, 41]. Recent efforts [73, 55, 12, 78, 11, 18, 82] have explored LLM-aided Hardware Description Language (HDL) generation and debugging using prompt engineering [20, 86, 57, 76], retrieval-augmented generation [46, 61], and supervised fine-tuning [36, 87, 83]. Besides HDL generation, some works have also investigated complex software designs. The learning selection method LAIL [40] and the self-planning code generation technique [30] are both enhancements in these areas. A two-stage code generation method [89] has demonstrated improved efficiency in generating code under zero-shot conditions. The LCG model [42], which adopts a multi-LLM agent cooperation mechanism, integrates seamlessly into traditional software development processes, significantly elevating the quality of generated code. In addition, study like UVLLM [27] has explored the potential of LLM in the generation of reference models. However, implementations rely on the relatively lower-level language SystemVerilog, which restricts the scale of the generated models. Despite these advancements, the potential of LLMs to generate reference models for large-scale hardware designs remains underexplored, with notable limitations in both the token/context memory limits for complex designs and limited debugging capabilities of LLMs.
3 Architecture & Workflow
Given that complex hardware designs consist of multiple functionally decoupled modules, ChatModel employs a “divide-and-conquer” approach to decompose intricate design tasks into manageable module-level subtasks, effectively addressing the aforementioned challenges presented in Section 1. Building on this strategy, ChatModel offers a fully automated solution for spec-to-model generation using LLMs, as shown in Figure 3. This framework consists of the “Design Specification Standardization Phase” and the “Hierarchical Agile Modeling Phase”. Each Agent is assigned a dedicated generation task and is equipped with a tailored chain of thought and template.
The transition from design specifications to reference model generation involves five key steps. (1) Module IRs Generation (Agents 1 & 2): Nonstandard input specifications are decomposed into multiple standardized module IRs based on the design structure, with each module IR corresponding to a specific generation subtask. (2) DFG Construction (Agent 3): For modules containing submodules, connections are established between parent module ports and internal submodule ports, effectively constructing the data flow graph. (3) Adaptive Task Planning: Utilizing Algorithm 1, the sequence for module generation and verification is defined, ensuring an organized and efficient workflow. (4) Module-Level Model Generation (Agents 4–6): Based on the module IR and DFG (optional), Algorithm 2 invokes the agile model generator to produce the corresponding reference model for each subtask. (5) Model Verification and Debugging (Agent 7): A comprehensive set of external test cases is employed to verify both the syntax and functionality of each generated module. Any module that fail verification are automatically repaired. If any modules remain incomplete after this process, subsequent modules are generated and validated iteratively until the top-level module is finalized, thereby completing the entire system design. To clearly demonstrate the workflow of ChatModel, Figure 4 illustrates the automatic generation and validation process of a reference model based on specific design specifications. For more detailed case information, please refer to Appendix E.
Figure 3: The workflow of the ChatModel Platform. In the first phase, the input design specifications are standardized into a design Intermediate Representation (IR). In the second phase, Hierarchical Agile Modeling (HAM) is employed to generate and verify reference models from module-subsystem-system based on the design IR.
Figure 4: Examples of Reference Model Generation Flow. The design consists of a top-level module (Conv2D) and its two submodules (Multi and Adder). In Step 1, module-level IRs for Conv2D, Multi, and Adder are generated sequentially using a top-down approach. Conversely, the reference model is built by first generating and verifying the Adder and Multi submodules, then the Conv2D top module. The final reference model consists of these three files.
4 Design Specification Standardization
The content of a design specification can vary significantly with different authors and often becomes excessively verbose, which diminishes the ability of LLMs to accurately extract design-related information. To address these, we propose a design specification standardization flow that employs multiple agents to decompose design at the module level, extract implementation descriptions for each module, and construct inter-module connections. Through this process, the input design specifications are transformed into a Design Intermediate Representation (Design IR), which consists of multiple module IRs and a Design Architecture Graph (DAG).
The design specifications are first divided into multiple text segments through preprocessing, which serve as inputs for the subsequent module-level IR generator. Module IRs are generated in a top-down systematic manner, starting from the top module and sequentially producing each submodule. As depicted in Figure 3, when Agent 1 is tasked with IR generation, it first sequentially extracts port signals, implementation details, and other relevant information from the sliced specifications of the current module. This extracted information is then analyzed and refined, with irrelevant content filtered out and implementation methods optimized, resulting in the final module IR. Agent 2 performs equivalence checks between the generated module IR and the original design specification to assess whether any essential descriptive logic has been lost during the transformation process. For IRs that fail verification, detailed error reports are communicated to Agent 1, accompanied by targeted guidance to facilitate iterative refinement and accurate regeneration. By employing multiple rounds of reading slice content and self-checking, ChatModel effectively mitigates attention loss in LLMs when processing long texts.
After the successful verification of all module IRs, Agent 3 constructs the DAG by systematically analyzing the port data flow graph of parent and child modules. The DAG abstracts the hierarchical structure of hardware modules into a structured directed acyclic graph, effectively resolving interface alignment challenges inherent in modular design. It comprises multiple depth-2 subtrees, where each subtree delineates the port connections between a parent module and its child modules, providing a precise and structured reference for subsequent generation processes.
5 Hierarchical Agile Modeling
In this section, we introduce the four key components of Hierarchical Agile Modeling: Design Planning, Agile Model Generator, Model Verification, and Model Debugger.
5.1 Design Planning
Input: Design Architecture graph $D=\{d_{1},\dots,d_{n}\}$
Output: Optimized Task Sequence $T_{\text{Seq}}$
1 Function Build_Seq($D_{\text{tree}}$, node, visited):
2       if node in visited then
3             return $\varnothing$
4       end if
5      visited.add(node);
6       $C=\{c\mid(p,c)\in D_{\text{tree}},p=\text{node}\}$;
7       $C_{\text{norm}}=\text{NormalizeChildren}(C)$;
8       $seq=\varnothing$;
9       //Post-order Traversal;
[…]
11             seq = seq + Build_Seq ($D_{\text{tree}}$, c_node, visited);
12             seq = seq + [node];
13
14       end foreach
15      return seq;
16
17
18Function HAP($D$):
19       // Parent-Child Relation Extraction;
20       $h=\varnothing$;
21       foreach $d_{i}\in D$ do
22             $(p_{i},c_{i})=\text{ExtractParentChildPairs}(d_{i})$;
23             $D_{\text{tree}}=D_{\text{tree}}\cup\{(p_{i},c_{i})\}$;
24
25       end foreach
26      // Root Identification;
27       $T_{\text{root}}=\text{IdentifyRoot}(D_{\text{tree}})$;
28       // Sequence Construction;
29       $\text{visited}=\varnothing$;
30       $T_{\text{Seq}}=\textnormal{{Build\_Seq}}(D_{\text{root}},T_{\text{root}},\text{visited})$;
31       return $T_{\text{Seq}}$;
32
33
34
Algorithm 1 Hierarchical Adaptive Planning
To meet the needs of generating reference models of various scales and agile verification of models, this work proposes a Hierarchical Adaptive Planning (HAP) algorithm to replace the traditional LLM random design task planning strategy. With HAP, ChatModel can automatically construct and execute the optimal task sequence even when the hardware design structure and scale change.
As shown in Algorithm 1, this algorithm preprocesses the input DAG by sequentially extracting parent-child module pairs $(P_{i},C_{i})$ and storing them in the design tree $D_{\text{tree}}$. Next, HAP identifies the root node $T_{\text{root}}$ of $D_{\text{tree}}$ and uses a post-order traversal strategy to traverse it. The algorithm explores deeper layers from the starting node, backtracking when no further exploration is possible, and continues along other unvisited paths. Since hardware designs often instantiate multiple modules with different names but the same functionality, several similar subtrees that do not require redundant generation may appear in $D_{\text{tree}}$. Therefore, HAP optimizes the generation of task planning paths by normalizing node names to identify and bypass certain subtrees, thereby improving the efficiency of reference model generation. The search strategy of HAP ensures that dependent modules are ready before subsequent modules are generated, allowing the validation of the internal subsystems during the generation process rather than waiting for the entire design to be completed.
5.2 Agile Model Generator
The Agile Model Generator (AMG) is responsible for delivering a stable and high-quality reference model based on the input Module IR and DAG. Within AMG, three cooperative agents manage the refinement of generation tasks, SystemC code generation, and code completeness checks. This collaborative process ensures that the resulting reference models meet the specified requirements.
Input: Module IR $M_{\text{IR}}$, Module DAG $\text{d}_{i}$
Output: Reference model $Ref_{\text{Model}}$
1 Function AMG($M_{\text{IR}}$, $\text{d}_{i}$):
2       // Functional block partition – Agent 4;
3       $f_{\text{blocks}}\leftarrow\text{aggregate\_func\_blocks}(M_{\text{IR}})$;
4       $\text{r}_{\text{c}}\leftarrow\varnothing$;
5       if module is parent module then
6             $r_{\text{c}}\leftarrow\text{analyze\_interactions}(M_{\text{IR}},\text{d}_{i})$;
7
8       end if
9      // Model generation & verification – Agents 5& 6;
10       $\text{c}_{\text{result}}\leftarrow\varnothing$;
11       while $\text{c}_{\text{result}}$ is not pass  do
12             $\text{code}_{\text{h}}\leftarrow\text{gen\_header}(M_{\text{IR}},f_{\text{blocks}},\text{r}_{c},\text{c}_{\text{result}})$;
13             foreach $f_{\text{block}_{i}}\in f_{\text{blocks}}$ do
14                   $info_{\text{i}}\leftarrow\{M_{\text{IR}},f_{\text{block}_{i}},\text{code}_{\text{h}},\text{r}_{c},\text{c}_{\text{result}}\}$;
15                   $\text{code}_{\text{s}_{i}}\leftarrow\text{func\_impl}(info_{\text{i}})$;
16                   $\text{code}_{\text{s}}\leftarrow\text{code}_{\text{s}}\cup\text{code}_{\text{s}_{i}}$;
17
18             end foreach
19            $\text{c}_{\text{result}}\leftarrow\text{eq\_check}(\text{code}_{\text{h}},\text{code}_{\text{s}},M_{\text{IR}},\text{d}_{i})$;
20
21       end while
22      $\text{Ref}_{\text{Model}}\leftarrow\{\text{code}_{\text{h}},\text{code}_{\text{s}}\}$;
23       return $Ref_{\text{Model}}$;
24
25
Algorithm 2 Automated Model Generation
As shown in Algorithm 2, Agent 4 invokes the corresponding module IR $M_{\text{IR}}$ based on the current generation task and employs CoT reasoning to analyze each functional point in the $M_{\text{IR}}$. Similar functional points are grouped into the same functional block, forming a set of blocks denoted as $f_{\text{blocks}}$, where each block corresponds to a single round of generation dialogue. This multi-round strategy ensures that the LLM focuses on generating one functional block per round, thereby minimizing incomplete outputs and hallucination errors caused by token limitations in LLMs. For parent modules, as certain child-module port signals may directly interact with the parent’s internal logic, Agent 4 refines the DFG to explicitly capture these interactions and stores the results in $r_{\text{c}}$.
[…]
5.3 Model Verification & Debugger
The validation strategy of ChatModel for the reference model adheres to the task planning generated by HAP. Upon generation of the reference model for a specific module, user-provided test cases are utilized to verify the syntax and functional accuracy of the reference model. The model is deemed functionally equivalent to the hardware design only after successfully passing all tests. Subsequent subsystem or system-level models will incorporate these validated models and undergo comprehensive subsystem or system-level validation. This incremental modular validation approach employs a bottom-up strategy, facilitating concurrent validation during the development of internal subsystems. In addition, it effectively reduces validation complexity, simplifies error localization, and accelerates convergence toward an accurate model.
To enhance the debugging capabilities of LLMs, we document common errors and their solutions within the reference model and utilize Chroma [64] to construct a vector database. Agent 7 employs Retrieval-Augmented Generation (RAG) technology to access relevant data from this database, facilitating the resolution of syntactic and functional errors during validation. For syntactic errors, terminal output logs are used to retrieve pertinent information from documents or knowledge bases via RAG technology. This information, along with relevant code snippets, is input into the LLM for error analysis and code correction. In the case of functional errors, Agent 7 also extracts logs from the period surrounding the error occurrence to enable the LLM to analyze the causes of output discrepancies.
6 Evaluations
Table 1: Comparison of pass@5 performance between ChatModel and LLM-based methods across various benchmarks. SP represents the Syntax Pass Rate, and FP denotes the Functionality Pass Rate. The highest scores in each column are highlighted with red font on a green background.
Method VerilogEval RTLLM2.0 ModelEval Average
SP/FP/SP/FP/SP/FP/SP/FP/%
Zero-Shot [84] 68.21 42.15 56.73 22.39 46.91 19.36 59.62 31.72
Few-Shot [80] 75.34 48.07 66.31 24.24 49.23 22.32 65.65 36.03
CoT [81] 73.47 51.44 67.14 27.37 53.82 23.11 66.26 38.55
CoT+Few-Shot [34] 87.41 53.21 71.04 31.53 55.14 28.31 71.45 41.79
Multi-Agent [58] 87.87 61.56 73.17 34.54 58.38 31.14 76.18 47.53
Best-Choice [60] 97.43 57.28 70.82 33.31 54.61 28.37 79.58 44.23
ChatModel 99.32 91.20 94.74 83.29 89.36 81.17 95.44 86.74
In this section, we evaluate the capability of ChatModel in generating reference models for various designs across multiple dimensions. The results demonstrate that the framework exhibits exceptional performance and remarkable stability when processing diverse designs, effectively enhancing the capability and efficiency of LLMs in agile model generation and verification tasks.
6.1 Experimental setup
Benchmarks. To evaluate the effectiveness of our work, we employed 300 designs from the open-source benchmarks VerilogEval [45] and RTLLM2.0 [49], as well as our own ModelEval dataset. Both VerilogEval and RTLLM2.0 are relatively simple and primarily used to evaluate the generative capabilities of LLMs, yet their design specifications can still be applied to generating reference models. To more accurately evaluate the capability of ChatModel in generating complex reference models, we developed the ModelEval benchmark, which consists of 94 distinct designs that vary significantly in complexity. Each design includes comprehensive test cases to verify the functional correctness of the generated reference models. The model is considered successfully generated only if it passes these test cases.
Evaluation Metrics. Since LLMs are probabilistic, each case is evaluated multiple times to account for output variability, and the pass@k [13] metric is calculated. Unless otherwise specified, ChatModel uses GPT-4o as the foundational LLM for each agent, and the temperature of all LLMs is set to 0.3 to balance generation capability and randomness. The debugging iteration threshold in ChatModel is set at 6, as improvements beyond this point are difficult to observe. For more detailed information about the experiment setup, please see Appendix B.
6.2 Capability for Reference Model Generation
Table 1 compares the reference model generation capabilities of ChatModel with CoT, Few-Shot, Multi-Agent and other LLM-based methods across three benchmarks. ChatModel consistently achieved the highest scores on both syntax and functionality tests, with an average syntax pass rate (SP) of 95.44% and a functionality pass rate (FP) of 86.74%. Compared to the baseline FP of the LLM-based Method, ChatModel demonstrates a performance improvement ranging from 39.21% to 55.02%, significantly improving the ability of LLMs to generate accurate reference models. In Figure 5, we selected several representative examples from the three benchmarks, arranged in order of increasing complexity. As the complexity of the designs increases, the scores of the baseline methods decrease rapidly, whereas ChatModel consistently manages to complete these designs effectively. The figure demonstrates that ChatModel, compared to other LLM-based methods, generating complex designs of over 1800 lines and enhancing design capability by 9.18x.
Figure 5: Comparison of Pass@5 between ChatModel and LLM-based Methods for specific designs.
Figure 6: Performance evaluation across different data sizes and Pass@k metrics. (a) compares average Pass@5 across design scales (small, medium, large) for four methods; (b) and (c) present the average Pass@k (k=1,5) rates for SP and FP across five tests, accompanied by error bars.
According to the design scale classification criteria specified in Appendix B, we further divided the benchmarks into three categories based on the design scale: Small-Set, Medium-Set, and Large-Set. We evaluated the average Pass@5 performance of ChatModel and other LLM-based methods on these datasets to compare their generative capabilities on different design scales, as depicted in Figure 6 (a). Zero-Shot and other methods perform reasonably well on small-scale design tasks, but their effectiveness declines markedly with medium-scale and large-scale designs. In contrast, ChatModel consistently exhibits superior generative pass rates across datasets of all sizes. In comparison to the second highest Multi-Agent approach, ChatModel achieves enhancements of 1.37x, 2.68x, and 10.23x on the three test sets, respectively, highlighting its substantial advantage in handling complex design tasks.
Additionally, we performed five repeated tests on all benchmarks with various k values, as shown in Figure 6 (b) and (c). The results indicate that ChatModel not only enhances the ability of LLMs to generate reference models, but also significantly improves the stability and robustness of the outputs. Under the pass@1 condition, ChatModel achieves a 51.62% improvement in the FP rate compared to the Zero-Shot baseline, with variability controlled at 7.45%. This suggests that the majority of tests consistently pass on the first attempt. Moreover, as the value of k increases, the improvement in the false positive rate increases to 55.02%, while the variability decreases to 4.08%.
6.3 Ablation Experiment
To further clarify the contribution of each component to the improvement in the generation capacity of ChatModel, we performed ablation experiments on designs of varying complexity. The entire experiment used GPT-4o as the original reference for comparison, and we sequentially added the four core elements of ChatModel: design IR, Hierarchical Adaptive Planning (HAP), Agile Model Generator (AMG) and debugger, to observe their impact on generation performance.
Figure 7: Ablation Experiments: Evaluating the Impacts of Design IR, Hierarchical Adaptive Planning (HAP), Agile Model Generator (AMG), and Debugger in ChatModel
Figure 8: Pass@5 (left) and input tokens (right) for each iteration.
Figure 9: Time Cost Distribution
Figure 7 demonstrates that the incorporation of design IR significantly improves generative performance, with HAP further enhancing the capabilities of LLMs in handling large and complex designs. This improvement is attributed to the effective reduction of hallucinations in LLMs through standardization of the specification, while HAP further mitigates the constraints imposed by token limitations. For more intricate designs, AMG effectively addresses the challenges of generating extended submodules and aligning interfaces. Furthermore, the introduction of a Debugger agent further strengthens the debugging capabilities of ChatModel.
The evaluation results shown in Figure 9 indicate that the Debugger component of ChatModel is capable of quickly correcting errors in the reference model in two iterations. In contrast, other baseline methods generally require more iterations and are heavily influenced by the quality of the initial output. Additionally, complex designs result in a rapid increase in context tokens across multiple iterations, further impairing the error detection and correction capabilities of LLMs.
6.4 Agile Design and Verification Analysis
To evaluate the acceleration provided by ChatModel in reference model design and verification, we engaged three verification engineers with 2 to 3 years of experience who are familiar with SystemC. They independently developed and verified reference models for three designs of varying complexity and recorded their time costs. Engineer E1 utilized ChatModel for development, E2 employed GPT-4o for assistance, while E3 developed without the aid of LLMs. As illustrated in Figure 9, ChatModel significantly improved development efficiency compared to E3, achieving an average speedup of 5.90×, with a range from 3.94× to 6.27×.
7 Conclusion
In this work, we introduce ChatModel, an agile framework for the generation and verification of reference models for general hardware. This framework effectively addresses the challenges faced in reference model design using LLMs through the standardization of design specifications and a structured Hierarchical Agile Modeling (HAM) process. We have also developed a debugger agent to effectively detect and correct errors in the generated models. Compared to traditional LLM-assisted reference model generation, our approach achieved a 55.02% improvement in functional pass rate and expanded the generative scale of the reference model by 9.18X. In comparison to manually developed reference models, ChatModel demonstrates an enhancement in reference model generation and performance ranging from 3.94X to 6.27X. Our evaluation demonstrates robust capability of ChatModel in generating complex reference models and further explores the transformative potential of LLMs in assisting reference model generation.
Appendix
Contents
[…]
0
Appendix A Challenges in Reference Model Generation with LLMs
To explore the challenges that large language models (LLMs) encounter when generating reference models, we conducted a comparative study involving GPT-4o, Claude 3.5-Sonnet, Gemma-2, Llama-3, and Code-Llama. We selected ten open-source designs with varying levels of complexity as the test set. Each LLM was tasked with generating reference models based on identical input specifications. A reference model was considered functionally equivalent to the hardware design only if it successfully passed all built-in functional verification cases. In instances of errors or test failures, the corresponding error logs were provided to the LLM for debugging. We then recorded the number of generation and debugging iterations required to achieve successful functional verification. Each design task was repeated 20 times, with a maximum of six debugging rounds per test. We evaluated the generation capabilities of the LLMs using Pass@k, as defined in Formula 1, and assessed their debugging capabilities using $D_{C}$, as defined in Formula 2.
Pass@$k$ $:=\mathop{\mathbb{E}}_{\text{Problems}}\left[1-\frac{{\binom{n-c}{k}}}{\binom{n}{k}}\right]$ (1)
$D_{C}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{Err_{i}+1}$ (2)
Here, Pass@k [13] is a widely used metric for assessing the generative capabilities of LLMs. This metric involves generating n code samples for each task (where n is greater than k), and then randomly selecting k samples. The test is deemed successful if at least one of these k samples passes the unit tests. The variable c represents the total number of samples that successfully pass the unit tests. In the formula of $D_{C}$, N denotes the number of debugging iterations required during functional testing, while $Err_{i}$ indicates the number of errors encountered in each debugging iteration. A higher score according to this formula signifies superior performance of the LLM.
Figure 10: LLM capabilities in reference model generation (left) and debugging (right)
As shown in the Figure 10, GPT-4o exhibits stronger generation and debugging capabilities compared to other LLMs. For simpler designs, LLMs can quickly generate the corresponding reference models. However, as design complexity increases, the ability of LLMs to generate accurate reference models diminishes. When the RTL code exceeds 160 lines, most LLMs struggle to produce correct reference models within a limited number of iterations. We attribute this occurrence primarily to four challenges:
  • Context Window & Hallucinations: Complex hardware designs are typically accompanied by extensive specifications, making it challenging for LLMs to systematically align and analyze the global context due to their limited token capacities and context window sizes. As a result, these limitations can induce hallucinations and significant errors in the generated reference models.
  • Randomness: LLMs can produce different outputs even with the same input. This inherent randomness increases the likelihood of design errors, which can accumulate rapidly over multiple iterations. The absence of standardized descriptive norms, combined with the diverse content and structure of complex design specifications, further exacerbates the problem and significantly affects output quality.
  • Token Constraints: The limited token generation capacity of LLMs makes it difficult to encapsulate all necessary details for large-scale designs in a single output. Therefore, multi-turn dialogues or multi-agent approaches are necessary. However, these methods require an effective strategy to manage the entire generation process and ensure that the interfaces within the design remain consistent.
  • Limited Code Generation & Debugging Capabilities: Hardware design necessitates consideration of timing and parallel execution, whereas reference models are typically executed sequentially. LLMs must accurately understand the distinctions between these two methodologies to generate reference models effectively. Moreover, the intricate inter-module connection relationships inherent in hardware designs result in error-prone code generations and challenging verifications, further impacting overall system reliability.
Appendix B Benchmarks, LLM-based Methods and Metrics
B.1 Benchmarks
In this section, we provide a comprehensive overview of the benchmarks utilized in our evaluation. Owing to the limited availability of open-source benchmarks specifically designed for reference models of hardware design, we have adopted VerilogEval and RTLLM2.0 for our experiments. Although these benchmarks primarily aim to assess LLMs in generating hardware description languages (HDLs), the provided design specifications can also be applied to generate SystemC reference models. By utilizing EDA tools such as VCS [71], the generated models can be co-simulated with the supplied test platforms to verify their functional equivalence to the hardware designs. Furthermore, we developed a benchmark called ModelEval to more accurately evaluate LLM performance under complex design requirements.
VerilogEval [45] is a benchmark developed by NVIDIA to evaluate the performance of LLMs in generating Verilog code for hardware design and validation. It includes 156 diverse generation tasks, spanning from simple logic circuits to complex finite state machines. As shown in the Table 2, each design is provided with four files: interface, specification, reference design, and testbench.
Table 2: Description of the File Composition in VerilogEval. Specifications are employed as inputs for the generation of reference models by LLMs, while testbenches are utilized to validate the accuracy of these models.
Category Function Description Example
Interface Records the input and output port signals and bit widths of design
input  [15:0] a,b,
input  sel,
output [15:0] out
Specification Describes the logical function implementation of the current design Create a 2-1 multiplexer. When sel=0, choose a. When sel=1, choose b.
Reference Design Serves as a “golden sample” for comparison with the generated design output under the same input stimuli
#Port Define
#Logic Implementation
assign out = sel ? b : a;
Testbench Provides comprehensive test cases and records the comparison results of the ref design and the design under test(DUT) under the same stimuli
# Test cases Generation
a<=15’b0000111111001110;
# Compare Out
if (out_ref !== out_dut)
 \$display("Has mismatches...");
else
 \$display("No mismatches...");
RTLLM2.0 [49] introduced by the Hong Kong University of Science and Technology, serves as a benchmark to assess the ability of LLMs to generate HDLs from natural language. This benchmark includes 50 diverse functional designs, categorized into four types: arithmetic, memory, control, and miscellaneous, with each design containing all the information specified in Table 2.
As shown in Table 3, the arithmetic design features various adders, subtractors, multipliers, dividers, comparators, accumulators, and other specialized units. The memory design is dedicated to data storage and retrieval, including FIFO (first-in-first-out) and LIFO (last-in-first-out) buffers, various shifters, as well as ROM and RAM. The control design manages state and counting mechanisms, incorporating finite state machines (FSMs), sequence detectors, and various counters, which are essential for controlling operational flow and ensuring the sequential execution of logic within digital systems. Finally, the miscellaneous design includes signal generation, RISC-V components, frequency dividers, and other modules designed for specific applications, showcasing a broad range of functionalities.
Table 3: Description of Design Categories in RTLLM2.0
Category Description Num Specific Designs
Arithmetic Common designs for arithmetic operations 19 Adders, subtractors, multipliers, dividers, comparators, accumulators, float units
Memory Designs for data storage and manipulation 7 FIFO, LIFO, shifters, LFSR, RAM, ROM
Control Designs for control and sequence detection 6 FSMs, sequence detectors, counters
Miscellaneous Designs for signal processing or specific applications 18 Signal generators, RISC-V components, frequency dividers, application-specific designs
ModelEval is a specialized benchmark intended for evaluating the generative capabilities of LLMs concerning reference models. In contrast to VerilogEval and RTLLM2.0, ModelEval is more specifically tailored to assess reference model generation capabilities and features a higher level of design complexity. It includes 94 distinct designs, each provided with design specifications, a golden reference, and an associated testing platform. These designs cover a wide range of types, including digital signal processing, image scaling, 3D accelerators, communication buses, and components of RISCV, with varying degrees of complexity. The table presents information on some of these designs, with code lengths ranging from 20 to 1800 lines and the number of internal submodules ranging from 1 to 21.
Table 4: Design modules and their functions in ModelEval
Design Code Lines Submodule Num Function
Norm 45 1 Floating-point data normalization
ARB 83 1 Memory access arbitration
AES 90 5 Image signal controller
WB_Reg 124 1 Wishbone interface register file
AE 197 1 Automatic exposure control processing
Timer 200 3 APB-compatible timer
Driver 224 2 LCD driver module
FIFO 282 6 Asynchronous FIFO
Fmul 320 2 Floating-point multiplier
SPI 387 3 SPI bus module
CAM 406 1 Camera-to-AHB data interface
SPDIF 420 2 SPDIF transmitter for digital audio
Matrix 843 14 Performs matrix operations
GEO 1814 21 Geometry processing engine
In summary, our benchmarks comprising 300 design tasks of varying complexity, encompassing a wide range of design categories such as common computational units, processors, memory, and application-specific integrated circuits (ASICs). The dataset is further divided into three subsets based on design scale: the small-set includes designs with fewer than 100 lines, the medium-set consists of single-module designs with 100 to 400 lines or multi-module designs with fewer than 300 lines, and the large-set encompasses all larger-scale designs. This segmentation facilitates a more effective evaluation of the performance of LLMs across datasets of different scales.
B.2 LLM-based Methods
In this subsection, we introduce several methods based on LLMs used in our experiments.
Zero-Shot prompting: Zero-Shot prompting leverages the capabilities of a general LLM to address new tasks directly. In this approach, the LLM does not require specific task examples but instead relies on the extensive knowledge and reasoning abilities accumulated during its pre-training phase to generate outputs. Consequently, this method allows the LLM to demonstrate strong generalization and adaptability, effectively performing various tasks even in the absence of task-specific information.
Few-Shot prompting: Few-Shot prompting utilizes an LLM to execute specific tasks with only a small number of examples. This approach maximizes the use of the LLM’s pre-trained knowledge by providing a limited number of examples to guide the model in understanding and completing the task without additional specialized training. Typically, few-shot learning involves including task instructions and examples in the input, enabling the LLM to infer task requirements from limited context and generate appropriate outputs.
Chain-of-Thought (CoT): Chain-of-Thought prompting guides the LLM to generate a series of intermediate reasoning steps, thereby aiding in more effective problem-solving for complex issues. Unlike traditional LLMs that typically generate outputs directly from inputs, chain-of-thought decomposes problems into multiple manageable sub-steps, allowing the LLM to better understand and resolve issues. This step-by-step reasoning process can significantly improve the LLM’s accuracy and efficiency in tasks requiring deep analysis, such as logical reasoning and code generation.
CoT+Few-shot: CoT+Few-shot enhances the LLM’s accuracy and efficiency in handling complex problems by presenting a series of step-by-step reasoning examples, thereby improving the model’s understanding of the problem’s structure and logic.
Multi-agent: The multi-agent approach involves decomposing a complex task into multiple subtasks and employing several LLMs to collaborate in completing them. The current LLM treats other models as black boxes, focusing solely on the output content without considering their internal dialogue information. This method allows the LLM to concentrate on the current task requirements, thereby enhancing the accuracy and robustness of the results.
Best-Choice: Best-Choice is also a multi-model generation strategy. This method involves aggregating the prediction results from multiple LLMs and selecting the highest quality output for subsequent processing. Consequently, this approach effectively reduces potential biases or errors from individual models, thereby improving the overall system performance.
B.3 Metrics
In our experiments, we employ the Pass@k metric, as defined in Equation 1, to assess the capability of LLMs in generating reference models for diverse designs. To evaluate test sets comprising multiple designs, we determine the average pass rate using Equation 3, where $N_{\text{Sum}}$ denotes the aggregate number of designs within the current dataset.
[…]
Appendix C Supplementary Experiment
To further evaluate the impact of various large language models on ChatModel and other LLM-based methodologies, we selected Claude 3.5-Sonnet [10] and DeepSeek-R1 [24] for testing. We employed the single-agent approach (CoT+Few-Shot) and the multi-agent approach (Multi-Agent) from Table 1, which exhibited superior performance as benchmarks for comparison with ChatModel. The experimental results, presented in Table 5, indicate that ChatModel substantially improves functionality pass rates, with a 48.81% increase when using Claude 3.5-Sonnet and a 37.01% improvement with DeepSeek-R1, relative to the Zero-Shot baseline. These findings suggest that the framework consistently exhibits robust performance and significant enhancement when integrated with different LLMs.
Table 5: Comparison of pass@5 performance between ChatModel and LLM-based methods across various benchmarks and LLMs. SP represents the Syntax Pass Rate, and FP denotes the Functionality Pass Rate. The highest scores in each column are highlighted with red font on a green background.
LLM Method VerilogEval RTLLM2.0 ModelEval Average
SP/FP/SP/FP/SP/FP/SP/FP/%
Claude 3.5 -Sonnet Zero-Shot 67.31 45.72 63.58 24.61 38.37 23.72 57.62 35.31
CoT+Few-Shot 79.97 59.87 74.09 28.84 44.65 29.53 67.92 45.19
Multi-Agent 76.83 52.14 69.24 32.11 46.91 32.55 69.32 42.66
ChatModel 93.61 89.17 88.69 80.36 84.76 77.19 92.73 84.12
DeepSeek -R1 Zero-Shot 78.22 60.31 69.92 27.85 35.42 20.95 63.43 42.57
CoT+Few-Shot 74.18 57.32 72.43 31.50 39.76 26.28 63.10 43.29
Multi-Agent 82.53 62.78 73.51 33.11 44.52 31.71 69.12 48.10
ChatModel 87.54 84.91 80.26 78.32 76.99 71.39 83.02 79.58
Appendix D Case Study
The ChatModel framework adeptly plans and generates processes based on input design specifications, incrementally creating reference models for various modules and executing automated validation and self-correction. By employing techniques such as task refinement, chain of thought, and templates, ChatModel effectively addresses challenges encountered by LLMs in code generation, such as hallucinations and token limitations, etc. In this section, we employ the Conv2D design illustrated in Figure 4 as a case study to demonstrate the flexibility, stability, and efficacy of ChatModel. The architecture of this design is organized with a top-level module, Conv2D, which integrates two submodules, Adder and Multi. These submodules are employed by Conv2D to perform logical processing operations.
D.1 Design IR Generation for Conv2D
The first phase involves the Conv2D design specification standardization process, which is divided into two steps to convert non-standard design specifications into a design Intermediate Representation (IR). The design IR comprises three module IRs: Conv2D, Adder, and Multi, along with a Design Architecture Graph (DAG) that illustrates the connections between the submodules and the parent module.
Figure 11: The example of transforming Conv2D design specifications into design IR. This stage is divided into two processes: module-level IR generation and DAG construction.
Utilizing a top-down approach, ChatModel decomposes complex design specifications into multiple standardized module-level IRs. As depicted in Figure 11, ChatModel initially engages Agent 1 to extract information pertinent to the top-level Conv2D module through multiple rounds of dialogue and generates a standard Conv2D Module IR. Subsequently, Agent 2 verifies the generated IR with the original design specification to ensure the conversion process is equivalent. Any erroneous information is relayed by Agent 2 to Agent 1 for reference in regenerating the IR. Upon successful check, ChatModel determines the subsequent IR generation task based on the “Children Modules” within the Conv2D IR.
After completing the IRs for the Conv2D, Multi, and Adder modules, these IRs are submitted to Agent 3 to construct the inter-module DAG. Agent 3 primarily focuses on the connectivity between module ports, while Agent 4 further elucidates the interaction between the internal logic of submodules and parent modules. This methodology allows the LLM to concentrate on analyzing uniform types of connection relationships, thereby enhancing the accuracy of DAG construction. Since Conv2D is a straightforward design with a depth of 2, it is only necessary to establish port connections between Conv2D and its submodules, Multi and Adder. For more intricate designs, this process may need to be repeated multiple times to construct several depth-2 subtrees.
D.2 Reference Model Generation for Conv2D
Figure 12: The example of the Hierarchical Agile Modeling process. The reference models for the Adder and Multi modules are generated and verified sequentially. Subsequently, the reference model for the top-level Conv2D module is completed.
Figure 13: The example of syntax and functional error fixes for the reference model of Conv2D.
Following the phase of the design specifications standardization, the IR and DAG files are delivered to the design planning to implement the Hierarchical Agile Modeling (HAM) process. HAM employs a bottom-up construction approach as illustrated in Algorithm 1, progressively generating the reference model of Adder, Multi, and Conv2D.
During the generation of the Adder module, ChatModel initially engages Agent 4 for functional partitioning. For simple modules such as Adder, partitioning is generally unnecessary, but for complex modules exceeding 500 lines of code, such division becomes essential. The partition results, along with the Adder IR, are forwarded to Agent 5 for reference model generation. Through a series of dialogues, the complete SystemC code for “Adder.h” and “Adder.cpp” is ultimately produced. Agent 6 is responsible for conducting integrity checks to ensure that no ports or logical functions are omitted during the generation process. If the check fails, the check result is returned to Agent 5 for self-correction. During the reference model generation for the Conv2D parent module, the DAG of Conv2D is provided to Agent 4 to clarify the interactions between the ports of submodules and the internal logic of the parent module. Agent 4 will analyze how the signal ports labeled "Inter" interact with the parent module’s internal logic. The refined DAG is then passed to Agent 5, with subsequent processing consistent with other modules.
When the reference model for the Adder module is completed, it undergoes sequential syntax and functional validation. Syntax validation typically relies on the compiler, whereas functional validation depends on externally provided test cases. If any stage fails, the process moves to debugging. Otherwise, it continues with the generation of reference models for the Multi submodules. As depicted in Figure 13, Agent 7 is assigned to correct syntax and functional errors in the design. For syntax errors, Agent 7 utilizes Retrieval-Augmented Generation (RAG) to gather relevant information, inputting both the code and terminal error messages into the LLM for analysis and correction. If the simulation is correct, but functional errors remain, the terminal outputs the inputs, actual outputs, and expected outputs at each time point. Agent 7 extracts information from 100ns before and after the output discrepancy and inputs the module IR into the LLM for comparative analysis and correction. Through multiple iterations of automated generation, validation, and verification, ChatModel sequentially produces reference models for Adder, Multi, and Conv2D. Once the functional validation of the top-level Conv2D module is successful, the reference model for the entire design is complete. Ultimately, the user receives a reference model comprising files for three module classes. The Conv2D model acts as the top level, internally invoking the classes of the other two submodels.
Appendix E Limitations and Future Work
As ChatModel is a framework composed of multiple agents, its generation quality and stability are influenced to some extent by the LLMs utilized by these agents. While our experiments with advanced LLMs like GPT-4o have yielded promising results for complex designs, tasks remain challenging for less capable LLMs. Despite employing strategies such as task segmentation, self-verification, chain of thought, and offering output templates to mitigate the likelihood of hallucinations, ChatModel cannot completely eliminate them. This issue may lead to reference models failing syntax and functionality tests, thereby affecting the generation efficiency of ChatModel. The completeness of functional verification for the reference model is typically influenced by the test cases. Future work will focus on fine-tuning LLMs according to specific generation tasks to enhance agent task handling capabilities and reduce hallucination occurrences, ultimately enabling automatic generation of reference models for ultra-large-scale designs like out-of-order multiprocessors.